Grade 9Mechanics

a car of mass 1000 kg rests on a horizontal road initially.it is then pushed over 5 m to a speed of 1 m/s before the engine is started .if the average frictional force acting on the car is 500 newtons the total work done in moving the car before the driver starts the engine is started

sumanth
10 Years agoGrade 9

2 Answers

ram charan
10 Years ago
v=u+at
1=at
s=ut+1/2at^2
u=0m/s
s=1/2at^2
at=1m/s
5m=t/2
t=10s
a=0.1m/s^2
w=mas-500*0.1
1000*0.1*5 – 50J
work done =+450J
Bandana
7 Years ago
Case 1
The car is pushed to a speed of 1m/s over a distance of 5m.
    So, the car is accelerated.
KE = ½ × m × v^2
      = ½ × 1000 × 1^2
      = 500 J
Work done against friction = Fs
                                           = 500 × 5
                                           = 2500 J
Total work done = 2500 + 500
                          = 3000 J